# LAPTH-015/24 Spherical symmetric solutions of

### v2

@@ 3.3 Case 3
As in subcases 2a and 2b, we choose $\Lambda=0$ to avoid naked singularities for $r\to\infty$. The metric function $y$ is
@@
with $k_{2}>0$ or ($k_{2}=0$ and $k_{1}>0$) for the solution to be stationary at spacelike infinity. The central singularity $r=0$ is always present. In the vacuum case $q=0$, the spacetime is a black hole with a single horizon. In the electromagnetic case $q\neq 0$, it can correspond either to a black hole with two horizons, an extreme black hole, or a naked singularity, depending on the parameter values.
@@ 4 The inverse problem
A static spherically symmetric matter source is characterized by a diagonal energy-momentum tensor $T_{\mu\nu}$, which may be thought of as a static anisotropic fluid, with components
@@ 5 Time-dependent vacuum solutions
Let us now integrate again this equation as follows. Putting $a=b^{-1}$ and using the conservation law (5.3), equation (5.4) may be rewritten as
@@
which (assuming $\dot{b}\neq 0$) is first integrated by
@@
.$\dot{a}^{2}+V(a)=0,\quad V(a)=-\alpha a^{4}-\displaystyle\frac{\Lambda}{3}a^{2}+k.$ (5.11)
@@ 6 Discussion
more general than (5.1). At any rate, it is clear that the Birkhoff theorem is not valid for conformal Killing gravity. We have shown that the Schwarzschild metric is not the unique static spherically symmetric solution to this theory. And there is at least a three-parameter manifold of time-dependent spherically symmetric solutions.

### v3

@@ 3.3 Case 3
As in subcases 2a and 2b, we choose $\Lambda=0$ to avoid a curvature singularity for $r\to\infty$. The metric function $y$ is
@@
with $k_{2}>0$ or ($k_{2}=0$ and $k_{1}>0$) for the solution to be stationary at $r\to\infty$. The central singularity $r=0$ is always present. In the vacuum case $q=0$, the spacetime is a black hole with a single horizon. In the electromagnetic case $q\neq 0$, it can correspond either to a black hole with two horizons, an extreme black hole, or a naked singularity, depending on the parameter values.
@@ 4 The inverse problem
We will not attempt here to answer this question in all its generality, but restrict our investigation to the case of a static spherically symmetric matter source  [^1] The spherical symmetry of $T_{\mu\nu\rho}$ does not necessarily imply the spherical symmetry of $T_{\mu\nu}$. . This is characterized by a diagonal energy-momentum tensor $T_{\mu\nu}$, which may be thought of as a static anisotropic fluid, with components
@@ 5 Time-dependent vacuum solutions
A first solution of this equation is the static solution $\dot{a}=0$, provided the density $\rho$ and pressure $p$ are time-independent. Then, rescaling the radial coordinate $r$ so that $a=1$, the FLRW metric (5.1) is of the form (2.15) with $y(r)=1$, $c=1$, $d=-k$, corresponding to the Einstein static universe. In the sourceless case ($\rho=p=0$), equation (5.4) gives $d=-\Lambda/6$. Note that the identification of the cosmological constant with an integration constant depends on which differential equations are integrated, the space-dependent equations for the functions $\nu(r)$ and $\lambda(r)$ in (2.8), or the time-dependent equation (31) of [18].
Assuming now $\dot{a}\neq 0$, let us now integrate again this equation as follows. Putting $a=b^{-1}$ and using the conservation law (5.3), equation (5.4) may be rewritten as
@@
which is first integrated by
@@
$\dot{a}^{2}+V(a)=0,\quad V(a)=-\alpha a^{4}-\displaystyle\frac{\Lambda}{3}a^{2}+k.$ (5.11)
@@ 6 Discussion
more general than (5.1). At any rate, it is clear that the Birkhoff theorem is not valid for conformal Killing gravity. there being at least a three-parameter manifold of time-dependent spherically symmetric solutions.